A manufacturer of colored candies states that 13 ​% of the candies in a bag should be​ brown, 14 ​% ​yellow, 13 ​% ​red, 24 ​%​blue, 20 ​% ​orange, and 16 ​% green. A student randomly selected a bag of colored candies. He counted the number of candies of each color and obtained the results shown in the table. Test whether the bag of colored candies follows the distribution stated above at the alpha equals 0.05 level of significance.

(A) Compute the expected counts for each color

(B) what is the test statistic

(C) What is the P value

Color
Brown 63
Yellow 63
Red 55

Blue

Orange

Green

59

87

65
Answer #1

Here expected value = P(Color) * Total value of that color/Total value

Here is the expected value table with chi-square table

=

Color Observed value Proportion Expected value Chi-square
Brown 63 0.13 50.96 2.845
Yellow 63 0.14 54.88 1.201
Red 55 0.13 50.96 0.320
Blue 59 0.24 94.08 13.080
Orange 87 0.2 78.4 0.943
Green 65 0.16 62.72 0.083
Total 392 1 392 18.473

Chi-square = = 18.473

dF = n -1 = 6 -1 = 5

p - value = CHIDIST(18.473,5) = 0.0024

that p - value is less than 0.05 so we reject the null hypothesis and conclude that the candies doesn't follow the pattern as provided.